OpenStudy (anonymous):

a store sells a 5 lb box of detergent for $5.25 and a 10 lb box of detergent for $9.75 ` which size box has the lowest price per pound ?

OpenStudy (anonymous):

5.25/5 =1.05 9.75/10=.975 so second is cheaper

OpenStudy (anonymous):

i think da five pound one wat grade r u in

OpenStudy (anonymous):

oh & im in 7th

OpenStudy (anonymous):

cool im in 9th

OpenStudy (anonymous):

yeaup `

OpenStudy (anonymous):

remember price per ounce means \[\frac{\text{price}}{\text{ounce}}\]